OpenStudy (anonymous):

How to solve integral tan^3 x in terms of sin and cos

zepdrix (zepdrix):

Do you mean, you need to convert to sines and cosines BEFORE solving the integral? Or you just need the final answer to be sines cosines?

OpenStudy (anonymous):

the answer key does it in terms of sec and tan, but I did it using sin and cos and got it wrong, but i think it should work out too..so i think i 've made some mistakes ...and so I want to know if it is possible to solve it by converting tan into sin and cos

zepdrix (zepdrix):

Hmm, it's much easier to do with tan and secant :O having trouble doing it that way?

OpenStudy (anonymous):

didnt realize u can use sec and tan >.> i tried it for 1 min but thought couldnt solve using that so went on with sin and cos

zepdrix (zepdrix):

\[\large 1+\tan^2x=\sec^2x \qquad \rightarrow \qquad \tan^2x=\sec^2x-1\]Using this identity gives us, \[\int\limits \tan^3x dx \qquad =\qquad \int\limits \tan x(\tan^2x)dx\] \[ \int\limits\limits \tan x(\sec^2x-1) dx \qquad = \qquad \int\limits \tan x \sec^2x dx - \int\limits \tan x dx\]

OpenStudy (anonymous):

so basically when I see tan, always try solving it using tan and sec right

zepdrix (zepdrix):

Ummmmm

zepdrix (zepdrix):

Yes, unless the problem involves any other trig functions. Because that might mean you can simplify things down by converting everything to sines and cosines :D

zepdrix (zepdrix):

Surprisingly, the level of difficulty of these integrals often comes down to the exponent on the trig function, whether it's even or odd.

OpenStudy (anonymous):

yea, i guess if the question already have sin and cos present, then it might be more helpful to convert it to sin and cos

OpenStudy (anonymous):

you always change the odd one right

OpenStudy (anonymous):

so it will be (x^2)(x^2) (x) < the x becomes the derivative

zepdrix (zepdrix):

Yes very good c:

zepdrix (zepdrix):

Odd is the one we like ^^

OpenStudy (anonymous):

just wondering what if they are both even

zepdrix (zepdrix):

err actually, EVEN is the one we want for Tangent i guess :) Odd for sine and cosine.

OpenStudy (anonymous):

i usuall just pick anyone lol, it should work out ..@@

OpenStudy (anonymous):

why even for tangent?

OpenStudy (anonymous):

they have similar trig identities though tan^x + 1 = sec ^2 x cos^x + sin^x = 1

zepdrix (zepdrix):

They don't have similar derivatives though. tangent produces a square. Sines and cosines don't :O

OpenStudy (anonymous):

sin^2 x + cos^2 x = 1 < is this one right.. they are all involving squares..

zepdrix (zepdrix):

the derivative of sine does not produce cos^2 right? :D

OpenStudy (anonymous):

oh..u mean the derivative , true :D

OpenStudy (anonymous):

then i guess i will search for sec/tan if it pops up an even function. and i'm pretty sure my prof will make it like an undoable question and i will just get stuck ..

zepdrix (zepdrix):

From where we broke down the tangent into 2 integrals, you understand how to do the first one? It's a pretty straightforward U-sub.

OpenStudy (anonymous):

yes i do. Can u help me with another hard integral question? just need a guide

OpenStudy (anonymous):

usuall its always deciding what to start with is the hardest part..

zepdrix (zepdrix):

Yesss so true! :D

zepdrix (zepdrix):

bring it :O I'll try to help XD

OpenStudy (anonymous):

thanks! integral of 1/ x^3 sqrt (x^2- 1)

zepdrix (zepdrix):

All in denominator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think has to use trig sub in here?

zepdrix (zepdrix):

Hmm yah it looks like it :O

OpenStudy (anonymous):

if i set x= sec theta, i will get a random tan..

zepdrix (zepdrix):

random? XD lol

OpenStudy (anonymous):

and cant even cancel out the x^3..

OpenStudy (anonymous):

lol..not random haha, something that i dont want it to appear in my equation lol

zepdrix (zepdrix):

So what you'll do is, Let, x=secQ dx=secQ tanQ dQ (those are thetas +_+) And then just plug everything in, and after much cancellations, convert leftovers to sines and cosines if you need to :O It shouldn't be too bad hmm

zepdrix (zepdrix):

Imma try it on paper real quick :D to make sure I'm not talkin out my butt

OpenStudy (anonymous):

haha thanks!

zepdrix (zepdrix):

yah it worked out really nice :O Need to see some steps? :D

OpenStudy (anonymous):

yes plz. want ot know what to do with the x^3

OpenStudy (anonymous):

x^3 = sec^3x?

zepdrix (zepdrix):

Yes :O true story.

OpenStudy (anonymous):

then do u have 1/ sec^2 theta somewhere ?

zepdrix (zepdrix):

1/sec^2? Hmm

zepdrix (zepdrix):

\[\large x=\sec \theta \qquad dx=\sec \theta \cdot \tan \theta \;d \theta\] \[\int\limits\limits\limits \frac{(dx)}{x^3 \sqrt{x^2-1}} \quad \rightarrow \quad \int\limits\limits \frac{(\sec \theta \cdot \tan \theta d \theta)}{\sec^3\theta \sqrt{\sec^2\theta-1}}\]

OpenStudy (anonymous):

i got \[\frac{ 1 }{ 2 }\sec ^-1 x + 2\frac{ \sqrt{1-x} }{ x }\] ....

zepdrix (zepdrix):

for what part? plugging everything in?

zepdrix (zepdrix):

Or that's what your integral gave you? I'm confused :D

OpenStudy (anonymous):

yea..final ans..ok..just ignore me then, lol i'm wrong

zepdrix (zepdrix):

Oh sorry I'm a little behind :D Was making sure you got the setup. Yes that looks close to the right answer, hmm

zepdrix (zepdrix):

Or is that from the book? :D

OpenStudy (anonymous):

no..I did it!

OpenStudy (anonymous):

xD

zepdrix (zepdrix):

Ew you draw your triangles toward the left! You're one of those!! lolol

OpenStudy (anonymous):

LOLLL

zepdrix (zepdrix):

The only part I'm a little worried about - You didn't apply the half-angle formula correctly :O it should have changed the 1/2 (1+cos2Q)

OpenStudy (anonymous):

do u?

OpenStudy (anonymous):

oh opps ..@@

OpenStudy (anonymous):

ahh..did it too fast..

OpenStudy (anonymous):

sin 2 theta / 2 .. can i still bring out the 1/2

zepdrix (zepdrix):

you mean write it as (1/2)sin2Q ... ? :o

OpenStudy (anonymous):

cuz ur equation is now 1/2 (theta + sin 2 theta / 2) to bring out the 1/2, dont u need to times theta by 2

zepdrix (zepdrix):

From here you need to remember your Double Angle Formula for Sine :O

OpenStudy (anonymous):

1/2 ( 1 - cos2x)!

zepdrix (zepdrix):

Integrating should have given you something like this, \[\large \frac{1}{2}\left(\theta+\frac{1}{2}\sin2\theta\right)\]

OpenStudy (anonymous):

yes , i am here

OpenStudy (anonymous):

i think i should use 2sinxcosx here tho?

zepdrix (zepdrix):

Yessss

zepdrix (zepdrix):

That's the double angle formula silly :O

zepdrix (zepdrix):

You were thikning of the half angle formula, with the 1/2(1-cos2x)

OpenStudy (anonymous):

I have no idea what those formulas called...LOL it is pain enuff to remember the formulas @@ .we are not given any formulas on the exam no calculators..

zepdrix (zepdrix):

No calc? ooo nice :D I wish i could take a class like that haha

OpenStudy (anonymous):

............I wish u can write it for me :( this is my last exam.. and i'm really not prepared.. i still have 12 more ques to go before finishing the first package

zepdrix (zepdrix):

poor guy :U